Required information (The following information applies to the questions displayed below.) Part 1 of 2 On January 1, 2021, Sp . (The following information applies to the questions displayed below.] On January 1, 2021, Splash City issues $400,000 of 8% b2. Record the bond issue on January 1, 2021, and the first two semiannua 2021. (If no entry is required for a particular tran
Answer #1

Cash interest = 400000*8%*6/12 16000.0 Interest Expenses =367422*9%*6/12 16534 367956*9%*6/12 16558 Amortization Sechedule InDate Issue Entry 1. Record the Bond Issue General Journal Debit Credit 01-01-2021 Cash 367422 Discount on Bonds Payable 32578